what is the measure of an angle whose supplement is 15 degrees less than twice the measure of the angle?
Answer by checkley77(12844)

x+(x-15)=180
2x-15=180
2x=180+15
2x=195
x=195/2
x=97.5 ans for the larger angle.
97.5-15=82.5 ans. for the smaller angle.
Proof:
97.5+82.5=180
180=180